# Speak the Same Language: Global LiDAR Registration on BIM Using Pose Hough Transform

### v1

@@ Abstract
The construction and robotic sensing data originate from disparate sources and are associated with distinct frames of reference. The primary objective of this study is to align LiDAR point clouds with building information modeling (BIM) using a global point cloud registration approach, aimed at establishing a shared understanding between the two modalities, i.e., “speak the same language”. To achieve this, we design a cross-modality registration method, spanning from front end the back end. At the front end, we extract descriptors by identifying walls and capturing the intersected corners. Subsequently, for the back-end pose estimation, we employ the Hough transform for pose estimation and estimate multiple pose candidates. The final pose is verified by wall-pixel correlation. To evaluate the effectiveness of our method, we conducted real-world multi-session experiments in a large-scale university building, involving two different types of LiDAR sensors. We also report our findings and plan to make our collected dataset open-sourced.
@@
In order to effectively deploy this method, it is important to consider certain hyper-parameters such as the size of the map and the height threshold. These parameters play a crucial role in achieving accurate alignment results and should be carefully tuned. In addition, it is worth noting that there are certain limitations to our current approach. Firstly, the environment should have a well-defined structural layout for the method to work effectively. Additionally, the current version of the method may have time constraints, and efforts should be made to reduce the computational time required for alignment. Future research and development should focus on refining the method to enhance its performance in these aspects.
@@ I Introduction
Three-dimensional (3D) point cloud data provides accurate modeling of the indoor environment. Its application extends across various phases in the construction, encompassing typical uses such as geometry quality inspection and real-time visualization [1]. Traditionally, obtaining precise point clouds often relied on stationary laser scanners. In the last ten years, mobile light detection and ranging (LiDAR) mapping has emerged as a well-established technique within the robotics community [2], like centimeter-level simultaneous localization and mapping (SLAM). This technique enables the fast acquisition of 3D point clouds using mobile robotic platforms, boosting the 3D point cloud-based applications in the construction industry.
@@
The utilization of 3D point clouds commonly involves the as-designed building data, e.g., visualizing elements on building information modeling (BIM) with augmented reality (AR) techniques or comparing LiDAR to BIM for quality management. Notably, a rigid transformation exists between the as-built source frame and the as-designed reference frame. Thus, frame alignment is indispensable for aligning the distinct sourced data within the same spatial space for subsequent analysis. In commercial software, users manually set the transformation to align the data, and manual tuning is necessary to ensure alignment quality. Regarding long-term robotic operations, external infrastructures are typically required to calibrate the different frames, like employing tags and markers in the building. However, these processes entail much human labor and additional costs, restricting the adoption of LiDAR point clouds in the construction industry, especially in large-scale environments.
@@
In this study, we propose an automatic alignment approach for LiDAR point clouds and BIM using a global point cloud registration approach, without the requirements of external infrastructures or manual verification. Compared to conventional point cloud registration tasks, we face three main challenges: 1) the modality of LiDAR point clouds and BIM is naturally different, which introduces difficulties in front-end feature extraction and data association; 2) indoor environments commonly contain repetitive patterns, leading to potential ambiguity for pose estimation; 3) the deviation of as-built and as-designed make the data association hard to obtain. To address these challenges, our method encompasses both the front end and back end. We employ lines as the fundamental representations of BIM and LiDAR data, enabling the generation of corner points and point-level correspondence estimation (data association). Subsequently, multiple pose candidates are obtained through voting under the Hough transform scheme. These pose candidates undergo verification using wall-pixel correlation to obtain the final pose estimation. We conduct experiments using real-world data from a large-scale university building. The dataset includes multi-session LiDAR point clouds collected from mobile mapping platforms with two different LiDAR sensor types.
@@
Overall, this study focuses on registering LiDAR data to BIM globally, and we summarize the contributions as follows,
@@
  • We design and develop a complete registration framework that could align LiDAR point cloud with BIM automatically and globally.
@@
  • A comprehensive descriptor is designed for cross-modality data association, which encodes the basic lines and corners in structural environments.
  • The pose estimation is achieved by Hough transform in parameter space and verified by wall-pixel correlation to guarantee robustness and accuracy.
@@
  • We conduct experiments in the real world using two different LiDAR sensors in a large-scale university building. We plan to open-source the LiBIM-UST dataset to benefit the community.
@@ II-A Robots Meet BIM
BIM serves as a digital representation of buildings and other physical assets in the architectural, engineering, and construction (AEC) industry. BIM generally encompasses comprehensive 3D spatial information, making it suitable for aligning robotic sensing data with the BIM model. This alignment process can effectively reveal deviations between the as-built and as-designed states, thereby benefiting the AEC industry. Conversely, BIM could also serve as a spatial map that facilitates robot localization and planning, eliminating the need for traditional mapping approaches. By leveraging the information provided by BIM, robots can navigate and operate within the environment without relying on explicit mapping procedures. This mapping-free characteristic of BIM enhances robot navigation capabilities. The subsequent paragraphs will delve into the relevant works conducted in these two aspects, namely aligning robot sensings to BIM for deviation analysis and utilizing BIM as a spatial map for robot navigation.
@@
In the robotics community, an existing recognition is that for long-term robots operating in stable environments, the mapping process of SLAM can be redundant as the generated map remains largely unchanged over time. As an alternative, utilizing Building Information Modeling (BIM) or low-level floorplans has emerged as a potential solution for long-term robot navigation. One notable approach is FloorPlanNet, proposed by Feng et al. [6], which segments rooms from point cloud data and aligns them with a floor plan. The core architecture of FloorPlanNet is based on a graph neural network for effective feature learning. Similarly, Zimmerman et al. [7] introduce a semantic visibility model integrated into a particle filter, enhancing long-term robot indoor localization. These works leverage floorplans as 2D maps, while BIM is a good choice for providing 3D point cloud maps. Yin et al. [8] generated a semantic-metric point cloud map from a BIM and designed a semantic-aided iterative closest point (ICP) algorithm for robot localization. This method effectively localizes a robot equipped with a moving sensor within the BIM environment. While the aforementioned works focus on point-level operations, there is also a promising direction in building instance-level robot navigation. In the study by Shaheer et al. [9], a novel map called S-Graph is designed based on BIM. The S-Graph consists of three-layered hierarchical representations, and a particle filter-based approach is employed to support robot navigation at the instance level.
@@
Point cloud registration, like iterative closest point (ICP) [10] and normal distributions transform (NDT) [11], focuses on local point-level alignment with a good initial guess. Global point cloud registration aims at solving pose estimation from scratch [12]. The global point cloud registration can be categorized into two primary models: correspondence-free[13] and correspondence-based. This study is in a correspondence-based manner. We mainly focus on point descriptors to provide correspondences and recent advances in this task.
@@
Figure 1: Schematic of our proposed method. The process begins with accumulating LiDAR scans as submaps (Section IV-A1). Feature extraction from the query submap, including 3D wall segmentation, 2D line detection, and corner detection, results in lines and corners in a LiDAR submap (Section IV-B). Feature extraction from Building Information Modeling (BIM) involves decomposition into subBIMs and extraction of lines and corners in the BIM data (Section IV-A2). These features are then used to compute triangle descriptors, which are stored in hash tables for efficient matching (Section IV-C). The back-end involves a pose Hough transform and voting procedure to identify potential transformation candidates (Section V-A). Finally, wall-pixel correlation verification is introduced to determine the optimal transformation (Section V-B).
@@ III System Overview
The primary objective of this paper is to achieve robust and global registration between Building Information Modeling (BIM) and Light Detection and Ranging (LiDAR) scans. The fundamental challenge arises from the heterogeneous nature of these modalities, which exhibit significant differences in structure, density, and information content. To address this, we have designed a pipeline that constructs correspondences using common features identified in the front end, and performs pose (transformation) estimation based on these correspondences in the back end.
The architecture of our proposed pipeline is illustrated in Figure 1. The process begins with data preprocessing, as outlined in Section IV-A, where LiDAR data are aggregated into submaps and the BIM model is decomposed into manageable subBIMs. This step is crucial to ensure consistency in density and size between both data sources. The subsequent extraction of common semantic features, specifically walls and corners, from these disparate data sources is discussed in Section IV-B. These features are then associated through the construction and matching of a novel triangle descriptor, as detailed in Section IV-C.
@@ IV Front-end: Walls, Corners and Descriptors
In this section, we establish correspondences between BIM and LiDAR submaps by extracting common semantic features such as walls and corners. The proposed pipeline consists of three main steps, as illustrated in Figure 1. First, sequential LiDAR point clouds are aggregated into a submap $\mathcal{S}$ to address the limited field of view issue, and the BIM is decomposed into several subBIMs ${}\{\mathcal{B}\}$ (Section IV-A). Second, walls and corners are detected and extracted from $\mathcal{S}$ and ${}\{\mathcal{B}\}$ to build correspondences, facilitate pose estimation, and perform verification (Section IV-B). Finally, corners are used to construct triangle descriptor dictionaries for precise point-level data associations (Section IV-C).
IV-A LiDAR Submap and SubBIM
The density and size of a single LiDAR scan and a BIM model often differ significantly due to modality differences. For effective construction monitoring, denser LiDAR data are preferable to capture detailed geometric information. Conversely, to reduce the complexity of the problem, preprocessing of LiDAR scans and BIM is necessary to align them in size and density. This involves accumulating sequential LiDAR scans to form the LiDAR submap $\mathcal{S}$, and decomposing the BIM into subBIMs ${}\{\mathcal{B}\}$.
IV-A1 Building LiDAR submap
@@
Given one LiDAR scan captured at the sensor frame $\mathcal{P}_{k}$, with $k$ representing the timestamp, we denote $\mathcal{P}_{1},\mathcal{P}_{2},\ldots,\mathcal{P}_{n}$ as a sequential of LiDAR scans with consecutive timestamps. A LiDAR submap could be obtained by processing sequential LiDAR scans:
$\mathcal{S}=\Gamma\left(\bigcup_{k=1}^{n}\mathbf{T_{k}}\mathcal{P}_{k},r_{v}\right)$ (1)
[…]
Figure 2: The decomposition process of BIM involves several steps. Initially, wall segments are extracted directly from the BIM using specialized software. Following this, a sliding window technique is applied to generate subBIMs. Within each subBIM, crucial features such as lines and corners are identified by utilizing the wall segments.
IV-A2 BIM decomposition
BIM data provides a digital representation of building elements, facilitating the direct extraction of semantic features such as walls using specialized tools or software, such as Revit [^1] Rhino.Inside.Revit at https://www.rhino3d.com/inside/revit/1.0/. After extracting these features, BIM decomposition is employed to manage the complex structure efficiently. This involves using a sliding window technique that moves along the $x$- and $y$-axes of the BIM floor plan at fixed intervals, with a predefined overlap ratio between each generated subBIM. The decomposition process is depicted in Figure 2.
@@ IV-B From 3D Walls to 2D Corners
In typical indoor environments, various semantic elements such as floors, columns, and walls exist [8]. Among these elements, the wall is a fundamental component of modern architecture. In this study, we propose selecting walls as the shared representations for LiDAR and BIM data. However, accurately modeling walls from LiDAR data using traditional point cloud processing techniques can be time-consuming. Additionally, learning-based wall recognition methods often require additional data to support training. These factors pose challenges in achieving large-scale global registration, especially on resource-constrained platforms.
On the other hand, the inertial measurement unit (IMU) is a gold standard for modern robotic sensing platforms. IMU can provide the gravity direction of the mapping platform, hence, the 6-degree of freedom (DoF) registration problem can be reduced to 3-DoF in the case of a planar floor. Thus, our basic idea is to project the 3D walls to 2D lines and extract corner points as features, thus generating descriptors for the downstream data association and 3-DoF pose estimation ($x-y$ coordinates and the yaw angle).
We first elaborate on the feature extraction for the LiDAR submap, as illustrated in Figure 3. With the built LiDAR submap, we perform plane detection on a single query submap to extract walls and filter out irrelevant components, notably the ground. Then the 3D walls are converted into 2D corners for the following descriptor computation and transformation estimation. Specifically, the entire point cloud is divided into voxels with predetermined dimensions $s_{v}$. Each voxel contains a collection of points $\mathbf{p}_{i}$, where $i=1,\ldots,N$. Next, we calculate the covariance matrix $\Sigma$ for the points within each voxel:
$\overline{\mathbf{p}}=\frac{1}{N}\sum_{i=1}^{N}\mathbf{p}_{i};\quad\boldsymbol{\Sigma}=\frac{1}{N}\sum_{i=1}^{N}\left(\mathbf{p}_{i}-\overline{\mathbf{p}}\right)\left(\mathbf{p}_{i}-\overline{\mathbf{p}}\right)^{T}$ (2)
To determine the content of each voxel, we analyze the ratio of the smallest eigenvalue $\lambda_{3}$ to the second smallest eigenvalue $\lambda_{2}$ of the covariance matrix $\Sigma$. If the ratio exceeds a predefined threshold, the voxel is considered to belong to a plane. To further consolidate plane voxels into coherent plane segments, we employ a region-growing technique. The criteria for merging voxels into plane segments are described in detail in our previous G3Reg by Qiao et al. [25].
Once plane segmentation is completed, we filter out ground points and project the remaining vertical wall points onto the floor plane. 2D corners are obtained by applying the following procedure:
@@
    We project wall points onto a 2D image at pixel scale $s_{\mathrm{I}}$ and apply a line segment detector to identify potential wall structures. We utilize the LSD algorithm by [26] for this purpose.
@@
    Detected line segments are then grouped based on their orientation. Segments whose lengths are below $L_{\mathrm{min}}$ will be discarded.
@@
    Within each group, the classical DBSCAN algorithm [27] is employed, using the distances between line segment endpoints to form clusters that likely correspond to individual walls.
  4. 4.
    We merge close and parallel line segments within the same cluster and refit them into a single line segment to accurately represent each wall.
  5. 5.
    Finally, these line segments are extended by a predefined distance, and their intersections are computed to identify corner points. These corners are crucial in subsequent data association and pose estimation.
Following a similar pipeline used for the LiDAR submap, we estimate corners of the subBIMs from extracted walls. This consistent approach across LiDAR submaps and subBIMs facilitates subsequent data association and back-end pose estimation.
Figure 3: The workflow for extracting triangle descriptors from LiDAR submaps begins with a Dense LiDAR Submap. The process involves plane segmentation to identify 3D walls. These segmented walls are then projected onto the 2D ground plane, allowing for the detection of lines. Subsequently, corner points are extracted from the detected lines through the intersection of adjacent lines. Finally, triangle descriptors are computed based on the corner triplets within the submap.

### v2

@@ Abstract
Light detection and ranging (LiDAR) point clouds and building information modeling (BIM) represent two distinct data modalities in the fields of robot perception and construction. These modalities originate from different sources and are associated with unique reference frames. The primary goal of this study is to align these modalities within a shared reference frame using a global registration approach, effectively enabling them to “speak the same language”. To achieve this, we propose a cross-modality registration method, spanning from the front end to the back end. At the front end, we extract triangle descriptors by identifying walls and intersected corners, enabling the matching of corner triplets with a complexity independent of the BIM’s size. For the back-end transformation estimation, we utilize the Hough transform to map the matched triplets to the transformation space and introduce a hierarchical voting mechanism to hypothesize multiple pose candidates. The final transformation is then verified using our designed occupancy-aware scoring method. To assess the effectiveness of our approach, we conducted real-world multi-session experiments in a large-scale university building, employing two different types of LiDAR sensors. We make the collected datasets and codes publicly available to benefit the community.
@@
Our proposed registration method leverages walls and corners as shared features between LiDAR and BIM data, making it particularly well-suited for scenarios with well-defined structural layouts. Accumulating a larger LiDAR submap provides richer structural information, which further aids in achieving accurate alignment. To optimize computational efficiency, we recommend constructing the descriptor database offline and loading it during runtime, enabling a theoretical retrieval complexity of $O(1)$. Despite its advantages, our approach has certain limitations. First, it primarily focuses on planar structures, which limits its effectiveness in utilizing non-planar features. Second, the method may underperform in cases where significant deviations exist between the as-designed BIM and as-is LiDAR data. Lastly, in ambiguous scenarios, such as long corridors or similar layouts within the same or across different floors, our method may struggle to verify the correct transformation among candidates. To address these challenges, incorporating additional information, particularly semantic cues such as floor numbers, room numbers, and room types, could enhance its robustness and reliability .
@@
Supplementary Materials
Video is available online.
The LiBIM-UST dataset of this study has been upgraded to the SLABIM dataset, available at https://github.com/HKUST-Aerial-Robotics/SLABIM.
@@ I Introduction
Three-dimensional (3D) point cloud data provides accurate modeling of the indoor environment. Its application extends across various phases in the construction, encompassing typical uses such as geometry quality inspection and real-time visualization [1]. Traditionally, obtaining precise point clouds often relied on stationary laser scanners. In the last ten years, mobile LiDAR mapping has emerged as a well-established technique within the robotics community [2], like centimeter-level simultaneous localization and mapping (SLAM). This approach facilitates the rapid acquisition of 3D point clouds using mobile robotic platforms, thus advancing 3D point cloud-based applications in both robotics and construction.
@@
In the architectural, engineering, and construction (AEC) industry, the utilization of 3D point clouds frequently involves the as-designed building data, e.g., visualizing elements of BIM with augmented reality (AR) techniques or comparing LiDAR to BIM for quality management. Notably, a rigid transformation inherently exists between the as-built source frame and the as-designed reference frame, making frame alignment essential for integrating distinct data sources into a unified spatial space for subsequent analysis. In other words, downstream comparisons and analyses involving BIM cannot be performed without proper frame alignment. In commercial software, users manually set the transformation to align the data, and manual tuning is necessary to ensure alignment quality. Regarding long-term robotic operations, external infrastructures are typically required to calibrate the different frames, like employing tags and markers in the building. However, these processes entail much human labor and additional costs, restricting the adoption of LiDAR point clouds in the construction industry, especially in large-scale environments. This frame alignment challenge also arises for mobile robots during global localization, where the robot must align itself with a pre-built map, such as BIM, as an initial step [3].
@@
In this study, we propose an automatic alignment approach for LiDAR point clouds and BIM using a global point cloud registration approach, without the requirements of external infrastructures or manual verification. Global point cloud registration has been studied in the field of robotics and computer vision, such as LiDAR-only registration [4, 5], and indoor RGBD alignment [6]. Compared to these tasks, we face three main challenges: 1) the inherent difference in modality between LiDAR point clouds and BIM, which introduces significant difficulties in front-end feature extraction and data association; 2) repetitive patterns in indoor environments, which can result in ambiguity for pose estimation; 3) the deviation between as-built and as-designed data, making reliable data association challenging to achieve. To address these challenges, our contributions encompass both front-end and back-end processes. We employ walls as the fundamental representations of BIM and LiDAR data, enabling the extraction of corner points and point-level correspondences (data association). Multiple pose candidates are subsequently obtained through voting based on the Hough transform and a hierarchical voting scheme. These pose candidates are verified using our proposed occupancy-aware score, leading to the final pose estimation. We validate our approach through experiments conducted on real-world data collected from a large-scale university building. The dataset includes multi-session LiDAR point clouds collected from mobile mapping platforms with two different LiDAR sensor types. Beyond its application in construction, the proposed global registration method can also be incorporated into robot navigation systems that leverage BIM as a navigational map [7, 8].
@@
Overall, this study focuses on globally registering LiDAR data to BIM, and we summarize the contributions as follows:
@@
  • We design and develop a complete registration framework capable of automatically and globally aligning LiDAR point clouds with BIM in one shot.
@@
  • A comprehensive descriptor with $O(1)$ retrieval complexity is proposed for cross-modality data association, effectively encoding basic walls and corners in structural environments.
  • A Hough transform scheme with hierarchical voting is introduced to hypothesize multiple pose candidates, followed by verifying the optimal transformation using an occupancy-aware score, improving both robustness and accuracy.
@@
  • We conduct real-world experiments using two different LiDAR sensors in a large-scale university building. Additionally, we plan to open-source the LiBIM-UST dataset and codes to benefit the research community.
@@ II-A Robots Meet BIM
BIM serves as a digital representation of buildings and other physical assets in the AEC industry. BIM generally encompasses comprehensive 3D spatial information, making it suitable for aligning robotic sensing data with the BIM model. This alignment process can effectively reveal deviations between the as-built and as-designed states, thereby benefiting the AEC industry. Conversely, BIM could also serve as a spatial map that facilitates robot localization and planning, eliminating the need for traditional mapping approaches. By leveraging the information provided by BIM, robots can navigate and operate within the environment without relying on explicit mapping procedures. This mapping-free characteristic of BIM enhances robot navigation capabilities. The subsequent paragraphs will delve into the relevant works conducted in these two aspects, namely aligning robot sensings to BIM for deviation analysis and utilizing BIM as a spatial map for robot navigation.
@@
In the robotics community, an existing recognition is that for long-term robots operating in stable environments, the mapping process of SLAM can be redundant as the generated map remains largely unchanged over time. As an alternative, utilizing BIM or low-level floorplans has emerged as a potential solution for long-term robot navigation. One notable approach is FloorPlanNet, proposed by Feng et al. [12], which segments rooms from point cloud data and aligns them with a floor plan. The core architecture of FloorPlanNet is based on a graph neural network for effective feature learning. Similarly, Zimmerman et al. [13] introduce a semantic visibility model integrated into a particle filter, enhancing long-term robot indoor localization. These works leverage floorplans as 2D maps, while BIM is a good choice for providing 3D point cloud maps. Yin et al. [8] generated a semantic-metric point cloud map from a BIM and designed a semantic-aided iterative closest point (ICP) algorithm for robot localization. This method effectively localizes a robot equipped with a moving sensor within the BIM environment. While the aforementioned works focus on point-level operations, there is also a promising direction in building instance-level robot navigation. In the study by Shaheer et al. [14], a novel map called S-Graph is designed based on BIM. The S-Graph consists of three-layered hierarchical representations, and a particle filter-based approach is employed to support robot navigation at the instance level.
@@
Point cloud registration, like ICP [15] and normal distributions transform (NDT) [16], focuses on local point-level alignment with a good initial guess. Global point cloud registration aims at solving pose estimation from scratch [3]. The global point cloud registration can be categorized into two primary models: correspondence-free[17] and correspondence-based. This study is in a correspondence-based manner. We mainly focus on point descriptors to provide correspondences and recent advances in this task.
@@
Figure 1: The pipeline of our proposed method begins with the accumulation of LiDAR scans to build submaps using LiDAR odometry (Section IV-A1). From these submaps, wall, corner, and ground features are extracted (Section IV-B) to construct triplets and generate triangle descriptors (Section IV-C). For BIM preprocessing, specialized software tools are employed to extract wall and corner features (Section IV-A2). The triangle descriptors are efficiently matched using a hash algorithm, enabling one-shot registration across the entire BIM. Based on the matched triplets, transformation candidates are generated, and a pose Hough transform combined with a hierarchical voting procedure is applied to identify the most plausible transformation candidates (Section V-A). Finally, an occupancy-aware confidence score is introduced to determine the optimal transformation (Section V-B).
@@ III System Overview
The primary objective of this paper is to achieve robust and global registration between BIM and LiDAR scans. The core challenge arises from the inherent modality difference, which manifests in variations in structure, density, and information content. To tackle this, we have developed a pipeline that first constructs correspondences by identifying common features in the front end and then performs pose (transformation) estimation based on these correspondences in the back end.
The architecture of our proposed pipeline is illustrated in Figure 1. The process begins with data preprocessing, as outlined in Section IV-A, where consecutive LiDAR scans are accumulated into LiDAR submaps using the relative pose estimates from the LiDAR Odometry. The subsequent extraction of common semantic features, specifically walls and corners, from the submaps and BIM is discussed in Section IV-B. These features are then associated through the construction and matching of a novel triangle descriptor, as detailed in Section IV-C.
@@ IV Front-end: Walls, Corners and Descriptors
In this section, we establish correspondences between BIM and LiDAR submaps by extracting common features: walls and corners. The proposed pipeline consists of three main steps, as illustrated in Figure 1. First, sequential LiDAR point clouds are accumulated into a submap $\mathcal{S}$ to improve the field of view, and the BIM is processed for element extraction ${}\{\mathcal{B}\}$ (Section IV-A). Second, walls and corners are detected and extracted in both $\mathcal{S}$ and ${}\{\mathcal{B}\}$ (Section IV-B). Finally, corners are used to construct triangle descriptor dictionaries for point-level data associations (Section IV-C).
IV-A LiDAR Submaps and BIM
IV-A1 Building LiDAR Submaps
The proposed approach for global LiDAR registration on BIM utilizes LiDAR submaps (also referred to as local maps). The preference for submaps over individual scans is based on two key reasons. First, submaps provide a more uniform point density than a single LiDAR scan, enhancing the accuracy of feature extraction for elements, as detailed in Section IV-B. Second, submaps cover a larger area compared to single scans, reducing ambiguity caused by limited viewpoint—particularly in indoor environments where views are frequently occluded—and thereby increasing the robustness of the registration process.
@@ IV-A1 Building LiDAR Submaps
Given one LiDAR scan captured at the sensor frame $\mathcal{P}_{k}$, with $k$ representing the timestamp, we denote $\mathcal{P}_{1},\mathcal{P}_{2},\ldots,\mathcal{P}_{n}$ as a sequential of LiDAR scans with consecutive timestamps. A LiDAR submap could be obtained by accumulating sequential scans with their relative pose transformations:
$\mathcal{S}=\Gamma\left(\bigcup_{k=1}^{n}\mathbf{T_{k}}\mathcal{P}_{k},r_{v}\right),$ (1)
where $\bigcup$ denotes the union of the point sets from each LiDAR scan; $\mathbf{T_{k}}$ denotes the transformation from the body frame to the world frame at each timestep, which is derived from LiDAR odometry [29]; $\Gamma(\mathcal{S},r)$ is the voxelization downsamling function, where $r$ represents the voxel size. The parameter $n$ is the minimum number of scans required such that the traversed distance exceeds $d_{s}$ meters.
IV-A2 BIM Preprocessing
We design and implement a pipeline to pre-process the BIM, focusing on extracting wall elements. The BIM model is first loaded in Revit [^1] Revit at https://www.autodesk.com/products/revit/overview/. Then we retain only wall components and filter out other elements according to the category by using Rhino.Inside®.Revit [^2] Rhino.Inside.Revit at https://www.rhino3d.com/inside/revit/1.0/, which is a plugin that enables data exchange between Rhino [^3] Rhino at https://www.rhino3d.com/ and Revit. Walls are then directly extracted using the “Element Curve” scripting component within the plugin. For each floor, the associated walls are projected onto a bird’s-eye view to identify corners and triangular descriptors, as illustrated in Section IV-B and IV-C.
It is worth noting that the BIM model is not divided into sub-areas. Instead, we perform a one-shot registration between the LiDAR submap and the entire BIM model. Furthermore, the BIM preprocessing, such as corner extraction and descriptor generation, is conducted offline. Thus, during the registration stage, the precomputed descriptor database of the BIM is loaded to ensure efficient use.
Figure 2: The workflow for extracting triangle descriptors from LiDAR submaps begins with a dense LiDAR submap. The process starts with plane segmentation to identify 3D wall structures. These segmented walls are projected onto the 2D ground plane, where lines are detected. Corner points are then extracted from the intersections of adjacent lines. Finally, triangle descriptors are computed based on the corner triplets within the submap.
@@ IV-B From 3D Walls to 2D Corners
In typical indoor environments, various semantic elements such as floors, columns, and walls exist [8]. Among these elements, the wall is a fundamental component of modern architecture. We analyze the BIM model of our university building and find that walls have the highest proportion ($>$30%) and distribution density compared to other elements. In this study, we propose selecting walls and corners as the shared features for LiDAR and BIM data.
We have introduced the wall extraction in the previous BIM preprocessing. The following part mainly focuses on feature extraction for LiDAR submaps. Figure 2 illustrates the pipeline for feature extraction from the LiDAR submap. First, plane detection is performed on a dense LiDAR submap to identify potential walls and the ground. To segment the planes, the point cloud data is discretized into voxels using an Octree structure, with the coarsest resolution denoted as $s_{v}$. Each voxel contains a collection of points $\mathbf{p}_{i}$, where $i=1,\ldots,N$. Then, the covariance matrix $\boldsymbol{\Sigma}$ for the points within each voxel is calculated as follows:
$\overline{\mathbf{p}}=\frac{1}{N}\sum_{i=1}^{N}\mathbf{p}_{i},\quad\boldsymbol{\Sigma}=\frac{1}{N}\sum_{i=1}^{N}\left(\mathbf{p}_{i}-\overline{\mathbf{p}}\right)\left(\mathbf{p}_{i}-\overline{\mathbf{p}}\right)^{T}.$ (2)
To determine whether the points in a voxel belong to a plane, we analyze the eigenvalues of the covariance matrix $\boldsymbol{\Sigma}$. Specifically, the ratio of the second smallest eigenvalue $\lambda_{2}$ to the smallest eigenvalue $\lambda_{3}$ is computed. If this ratio exceeds a predefined threshold $\sigma_{\lambda}$, the voxel is considered to belong to a plane.
If the ratio does not meet the threshold, the voxel is further subdivided into four sub-voxels, and the plane identification process is repeated until each sub-voxel contains fewer than four points. Next, for voxels identified as belonging to a plane, neighboring planar voxels are merged if their normal directions are similar and their point-to-plane distances are small. This merging process enables the extraction of complete structures, such as walls or the ground.
Once plane segmentation is completed, the wall and ground planes are distinguished based on the angular difference between their normals and the gravity direction provided by widely used IMU. The process of extracting 2D corners from 3D walls is as follows:
@@
    Wall points are projected onto the ground plane and rasterized at a pixel scale of $s_{\mathrm{I}}$. A line segment detector based on Hough Transform is then applied to identify potential wall structures.
@@
    We discard 2D line segments when they are shorter than $L_{\mathrm{min}}$ to eliminate insignificant segments. The remaining 2D line segments are merged based on the distance between their endpoints and the similarity of their directions. The merged segments are then refitted using the original 2D points to accurately represent the wall geometry.
@@
    Finally, the refitted line segments are extended by a predefined distance to intersect and determine the corner points. To ensure the extracted corners are not too close to each other, they are refined using Non-Maximum Suppression (NMS).
Following a similar pipeline used for the LiDAR submap, we estimate corners of the BIM from the extracted walls. This consistent approach across LiDAR submaps and BIM facilitates subsequent data association and back-end pose estimation.